There are four questions in tnis assigniment. Tou are required to show your working f question in the blank space provided. 1. A 2.5 ml. syringe has an inner diameter of 5 mm a needle of inner diameter of 0.32 mm, and a plunger pad diameter (where you place your finger) of 1 cm. A nurse uses the syringe to inject medicine into a patient whose blood pressure is 120/80 mmllg. Assume the liquid is an ideal fluid. A) What is the minimum force the nurse needs to apply to the syringe? B) The nurse empties the syringe in 5 s. What is the flow speed of the medicine through the needle? d) Home F6 F7 F8 F9
Answer #1

A/

blood pressure

p = 120/80 mmHg = 15999 Pa

force against end of syringe

F = p*A = 15999Pa * π(0.00500m)²/4 = 0.314 N

B/

Q = V / t = 2.5mL / 5s = 0.5 mL/s = 0.5 cm³/s

Then velocity

v = Q / A = 0.5cm³/s / (π(0.032cm)²/4) = 622 cm/s

or 6.22 m/s